Approximate to the nearest inch,the length of a rectangle whose diagonal measures 25.0 inches and whose width measures 18.0 inches
Answer by nerdybill(7384)

Approximate to the nearest inch,the length of a rectangle whose diagonal measures 25.0 inches and whose width measures 18.0 inches
.
Apply Pythagorean Theorem:
Let x = length
then
x^2 + 18^2 = 25^2
x^2 = 25^2 - 18^2
x^2 = 625 - 324
x^2 = 301
x = 17.35
rounded to nearest inch:
x = 17 inches